OpenStudy (anonymous):

need to figure out to set up this problem? 2 less than 8 times an unknown # is -74, what is the unknown #

OpenStudy (anonymous):

I think you set it up like this but cant remember 2-8x=-74 2+2-8x=-74+2

OpenStudy (anonymous):

If it is 2 less than 8 times an unknown, then it would be 8x - 2 = -74. 2 - 8x would be 8x less than 2. Do you understand what I am saying?

OpenStudy (anonymous):

yes I understand now

OpenStudy (anonymous):

ok how about the quotient of two integers is 5 one of the integers is -2 find the other integer?

OpenStudy (anonymous):

normally I don't have issues with these but cant get it straight to save my life

OpenStudy (anonymous):

okay. So we know that \[\frac{ a }{ b} = 5\] and that either 'a' or 'b' is equal to -2. Now since we know that we are dealing strictly with integers, we can rule out all decimals. This means that -2 must be the denominator, because if you take the absolute value of an integer and divide it by any number you cannot have an answer that is greater than the integer on top. Following so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so you set it up like this (x/2)*(2)=5*2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

yes. Close. remember that we are plugging in -2 though and not 2.

OpenStudy (anonymous):

yea forgot to type -2

OpenStudy (anonymous):

so the answer I got was x=10

OpenStudy (anonymous):

5*-2=10

OpenStudy (anonymous):

close. a positive times a negative yields a negative answer remember (darn those pesky negative signs lol) so the correct answer would be -10 not 10

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so pretty much when it says for instance 6 less than 3times a unknown number you would put 3x-6 right?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

k I think I have it now thank you

OpenStudy (anonymous):

I add 6 to the answer like this 3x-6=21+6 cause you do the opposite on the other side of the equation

OpenStudy (anonymous):

I think you understand it. But be careful how you phrase it, if you simply say you "do the opposite on the other side of the equation" it sounds like you are saying that 3x-6 =21 AND 3x-6 =21 +6. The more accurate statement would be "whatever you do to one side, you have to do to the other side too."

OpenStudy (anonymous):

yea that's what I was trying to say

OpenStudy (anonymous):

Okay, thought so, but wanted to be sure. I find that when I start phrasing how I work out a math problem in an ambiguous way, I tend to start making silly mistakes because even though I knew what I should be doing, I inadvertently did something else.

OpenStudy (anonymous):

I have one more 4 more then 9 times an unknown #is 85 in the book its set up like this 4+9x=85

OpenStudy (anonymous):

Correct so far.

OpenStudy (anonymous):

so it would be 4-4+9x=85-4

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

good cause that's how I had it on my paper at first I thought it was wrong the way I was writing it

OpenStudy (anonymous):

thank you for the help gotta eat dinner